a. Derek can deposit $11,000 on each birthday beginning with his 26th and ending with his 65th. What will the rate on the retirement account need to be for him to have $3,000,000 in it when he retires on his 65th birthday?

b. Derek can deposit $11,157.00 on each birthday beginning with his 28.00th and ending with his 69.00th. What will the rate on the retirement account need to be for him to have $3,168,894.00 in it when he retires?
Answer #1

a. Number of deposits = 65 - 26 + 1 = 40

N = 40

PMT = -11,000

PV = 0

FV = 3,000,000

CPT I/Y

I/Y = 8.19499304

The rate on the retirement account = 8.19499304%

b. Number of deposits = 69 - 28 + 1 = 42

N = 42

PMT = -11,157

PV = 0

FV = 3,168,894

CPT I/Y

I/Y = 7.75372566

The rate on the retirement account = 7.75372566%

Can you please upvote? Thank You :-)